Write a quadratic equation that has roots -5/2 and 3. Could you please show me how to work this problem. Thank you
Answer by stanbon(75887)

Write a quadratic equation that has roots -5/2 and 3.
If x=-5/2 then 2x+5=0
If x=3 then x-3=0
Multiply these to get the quadratic:
(2x+5)(x-3)=0
y=2x^2-x-15
Cheers,
Stan H.